OpenStudy (anonymous):

Can someone help me with this question: Let W be the set of all vectors v = (v1,v2,v3) in R3 such that v1-2v2+v3 =0 Prove that W is a subspace of R3

OpenStudy (fwizbang):

Assume v satisfies the conditions you listed above and let c be a scalar. Does cv still satisfy the conditions? Now assume there is a second vector w that also satisfies the conditions. Does v+w satisfy the conditions? If the answer to both questions is yes, then you have a subspace.

OpenStudy (anonymous):

Definition: \[v_1 - 2v_2 = v_3 = 0: v_3 = 0, v_1 = 2v_2\]\[w -> (2v_2,v_2,0)\] Associative Condition: \[w + w' = (2v_2,v_2,0) + (2v'_2,v'_2,0) = (2(v_2+v'_2),(v_2+v'_2),0) = (2v''_2,v''_2,0)\] where \[v''_2 = v_2 + v'_2\] Existence of Identity: \[w + e = w\] \[(2v,v,0) + (2x,x,0) = (2v,v,0)\]\[=> (2(v + x),(v + x),0) = (2v,v,0)\]\[=> v + x = v\]\[=> x = 0\]\[e -> (0,0,0)\] Existence of Identity: \[(2v,v,0) + (2v',v',0) = (0,0,0)\]\[(2(v + v'),(v + v'),0) = (0,0,0)\]\[v + v' = 0\]\[v' = -v\]\[(2v,v,0)^{-1} = (-2v,-v,0)\] Therefore W is a group and since it is within V, it is a subgroup of V

OpenStudy (anonymous):

Thanks everyone for your help. It is appreciated.

OpenStudy (anonymous):

check 3 conditions 1> (0,0,0) is there in subspace 2> scalar multiplication is there in subspace 3> addition of any arbitrary vectors must be there in subspace